# FIAN/TD/12-2024

### v2

@@ Abstract
In this note we construct a formulation of $4d$ pure Yang–Mills theory in the unfolded language of higher-spin gravity. It represents a first-order formulation possessing manifest diffeomorphism- and gauge-invariance. We make use of an unfolding method that was recently put forward and used to unfold scalar electrodynamics. We also discuss features of various unfolding maps defined by the unfolded equations.
@@ Introduction
Symmetry considerations are among the most important guiding principles in constructing theories of fundamental interactions. Higher-spin gravity, representing a theory of interacting massless fields of all spins, sets an example. To a large extent, this theory is determined by an infinite-dimensional higher-spin gauge symmetry [1]. For a partial review of the literature on the subject see [2].
To keep the higher-spin symmetry under control is of crucial importance. Originally, higher-spin gravity was formulated in the form of Vasiliev equations [3, 4] which are classical equations of motion written in the so-called unfolded form. Unfolded equations are first-order differential equations on unfolded fields being exterior forms, which possess manifest gauge symmetry. Thus, the unfolded dynamics approach [5, 3, 4, 6, 7, 8] represents a classical first-order formalism providing manifest diffeomorphism- and gauge-invariance. It also allows for an effective control over degrees of freedom, since unfolded fields (whose spectrum is infinite in dynamical field theories) parameterize all d.o.f. of a theory under consideration. In particular, this makes the unfolded dynamics approach a promising tool for the study of $AdS/CFT$-correspondence and other dualities [9, 10, 11]. A way to quantize classical unfolded theories was put forward in [12].
Apart from higher-spin gravity models [3, 4, 13, 14, 15, 16], not that many examples of unfolded formulations of nonlinear theories are available [17, 12, 18], which limits the ability to explore the full potential of the approach. The reason behind this was a lack of a practical algorithm of unfolding, which was especially critical for nonlinear theories. Therefore, linear models were mainly studied [19, 20, 21, 22, 23]. In [18], a novel unfolding method was put forward, which allowed one to construct an unfolded formulation of $4d$ scalar electrodynamics. In brief, it consists in the initial postulation of the form of an unfolded master-field and the subsequent derivation of unfolded equations as identities for this master-field.
In this note, we apply the method of [18], slightly improving it, in order to construct an unfolded formulation of $4d$ on-shell pure Yang–Mills theory.
The note is organized as follows. First, we present basic definitions of the unfolded dynamics approach. Next, we construct unfolded Yang–Mills equations by fixing the form of the unfolded master-fields and using certain operator relations that we derive. Then we discuss the properties of the resulting unfolded system and comment on the unfolding maps it defines. In conclusion, we provide a brief overview of possible directions for further research.
@@
Within the unfolded dynamics approach [5, 3, 4, 6, 7, 8], a field theory is formulated in terms of first-order equations
@@
on unfolded fields $W^{A}(x)$, which are exterior forms on a space-time manifold $M^{d}$. Here $A$ stands for all indices of an unfolded field, $\mathrm{d}$ is the exterior differential on $M^{d}$ and $G^{A}(W)$ is built from exterior products of unfolded fields (the wedge symbol is omitted throughout the paper). For every $W^{A}$ there is one and only one unfolded equation (1).
There is a consistency condition for the system (1)
@@
which follows from $\mathrm{d}^{2}\equiv 0$ and restricts possible $G^{A}$. Considering this, unfolded equations (1) possess a manifest gauge symmetry [^1] Strictly speaking, this is true provided that consistency holds regardless of the dimension of $M^{d}$, which is the case for all known examples [8].
@@
so that every $(n>0)$-form field $W^{A}$ gives rise to a gauge symmetry with a $(n-1)$-form parameter $\varepsilon^{A}(x)$. 0-form fields are transformed only under gauge symmetries of 1-form fields via the second term in (3).
In dynamical field theories, a spectrum of unfolded fields is infinite, since they parameterize all d.o.f. Usually, the space of unfolded fields possesses some grading bounded from below, so that equations (1) express (perhaps, in a very complicated nonlinear way) higher-grade fields in terms of derivatives of the lower-grade ones. In addition, (1) may implicitly impose dynamical constraints on the lowest-grade fields, in which case the system is said to be on-shell.
Ultimately, an unfolded system (1) describes a theory of lowest-grade fields (also referred to as primaries), possibly subjected to some differential constraints (e.o.m. or whatever), while higher-grade fields form (infinite) towers of their covariant differential descendants. Importantly, an unfolded formulation is manifestly diffeomorphism- and gauge-invariant and provides the control over d.o.f. of the theory. All this makes unfolded dynamics approach a powerful tool for studying higher-spin gravity. But besides higher spins, these remarkable features can be useful for more standard field theories as well. Here our goal is to reformulate $4d$ pure Yang–Mills theory in the unfolded form (1), making use of the unfolding method proposed in [18].
@@
In $sl(2,\mathbb{C})$-spinor notations, $4d$ Yang–Mills equations together with Bianchi identities are
@@
where the (anti-)self-dual components of the strength tensor are
[…]
@@
so that
$[\mathrm{D}_{\alpha\dot{\alpha}},\mathrm{D}_{\beta\dot{\beta}}]=-i\epsilon_{\alpha\beta}\bar{F}_{\dot{\alpha}\dot{\beta}}-i\epsilon_{\dot{\alpha}\dot{\beta}}F_{\alpha\beta}.$ (7)
Antisymmetric spinor metric
@@
moves spinor indices up/down according to
@@
Indices of multispinors, denoted with the same letter, are either contracted or symmetrized, depending on their relative positions (the same holds for dotted indices)
@@
In order to unfold Yang–Mills theory, one has to introduce, on top of primaries $F_{\alpha\alpha}$ and $\bar{F}_{\dot{\alpha}\dot{\alpha}}$, the infinite towers of all their differential on-shell descendants. This can be conveniently performed by means of auxiliary commuting spinors [^2] In $4d$ Vasiliev equations [3, 4], these spinors become generators of an associative higher-spin gauge algebra through a certain non-commutative star product defined on them. $Y=(y^{\alpha},\bar{y}^{\dot{\alpha}})$, then the whole towers get packed into unfolded Yang–Mills master-fields, which we postulate to be of the form
@@
together with an infinite sequence of their fully symmetrized traceless covariant derivatives of all orders. This constitutes a set of all independent covariant descendants of the primary Yang–Mills tensor, since antisymmetrizations and contractions are determined by (4) and (7). By construction, unfolded master-fields (11) inherit the adjoint representation of the gauge algebra of primary tensors.
@@
where $\partial_{\alpha}$ and ¯$\bar{\partial}_{\dot{\alpha}}$ are $y^{\alpha}$- and $\bar{y}^{\dot{\alpha}}$-derivatives. From their definitions (11), master-fields obey
@@
which corresponds to helicities ±$\pm 1$ (in higher-spin gravity, strength tensors with $|N-\bar{N}|F=2sF$ describe spin-$s$ massless fields) and, further,
@@
so that the primary has 00-grade, while $n$-grade unfolded fields represent its covariant derivatives of $n$-th order, as follows from (11).
To formulate the unfolded equations (1), one needs to express the derivatives of the unfolded fields in algebraic terms. In our particular case, the task is to express $\mathrm{D}_{\alpha\dot{\beta}}F$ for $F$ (11) in terms of $Y$ and $\partial/\partial Y$ acting on $F$ and $\bar{F}$ (and the same for $\mathrm{D}_{\alpha\dot{\beta}}\bar{F}$). Before we start processing (11), let us work out some general operator relations, which simplify the analysis implied by the method of [18]. Combining a commutator formula
@@
with an Euler-operator representation of a homotopy integral
@@
one obtains
$[\hat{\mathrm{A}},e^{\hat{\mathrm{D}}}]=(\frac{1}{N_{\hat{\mathrm{D}}}}e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}})e^{\hat{\mathrm{D}}}=e^{\hat{\mathrm{D}}}(\frac{1}{N_{\hat{\mathrm{D}}}}e^{-\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{\hat{\mathrm{D}}}),$ (21)
where the inverse Euler operator is understood as (for $F(0)=0$)
[…]
Applying this to an operator of the form
@@
$[\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}}e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}.$ (24)
@@
$e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}=[\hat{\mathrm{A}},\hat{\mathrm{D}}]-\frac{1}{N_{\hat{\mathrm{D}}}-1}(e^{\hat{\mathrm{D}}}[[\hat{\mathrm{A}},\hat{\mathrm{D}}],\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}),$ (25)
@@
$[\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}(N_{\hat{\mathrm{D}}}-1)}e^{\hat{\mathrm{D}}}[\hat{\mathrm{D}},[\hat{\mathrm{A}},\hat{\mathrm{D}}]]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+[[\hat{\mathrm{A}},\hat{\mathrm{D}}],\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}.$ (26)
For convenience, we denote an "unfolding map" for an arbitrary function $\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)$ as
@@
Then for an arbitrary $\mathrm{C}(Y|x)$ one has from (26)
$\ll\partial_{\mu}\mathrm{C}\gg=(\partial_{\mu}-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}})\ll\mathrm{C}\gg+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\ll\mathrm{C}\gg],$ (29)
¯$\ll\bar{\partial}_{\dot{\mu}}\mathrm{C}\gg=(\bar{\partial}_{\dot{\mu}}-\mathrm{D}_{\alpha\dot{\mu}}y^{\alpha})\ll\mathrm{C}\gg+i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+1)(\bar{N}+2)}F,\ll\mathrm{C}\gg].$ (30)
From here on, all derivatives inside the angle brackets always act only on an expression within brackets and never differentiate unfolding exponents of (27). The square brackets stand for the commutator in the gauge Lie algebra and the Euler operators of (26) are expressed in terms of spinor Euler operators (14). Master-fields $F$ and $\bar{F}$ arise in (29)-(30) through (7) without assuming Yang–Mills equations (4).
Now let us get down to solving the problem. Direct application of (24) to $\mathrm{D}_{\mu\dot{\mu}}F(Y|x)$ gives
@@
The task is to get rid of all covariant derivatives and angle brackets on the r.h.s. by re-expressing them in terms of unfolded master-fields $F$ and $\bar{F}$ acted on by $Y$’s and $\frac{\partial}{\partial Y}$’s. To accomplish this, we have the following tools at our disposal: Yang–Mills equations (4) together with (7), relations (29) and (30), Schouten identities for spinors and Jacobi identity of the gauge Lie algebra.
@@
So, one needs to process $\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F$. Contracting (32) with $\bar{y}^{\dot{\mu}}$ gives
@@
¯$\ll\bar{\partial}_{\dot{\mu}}\bar{F}_{\dot{\alpha}\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg=\frac{2}{\bar{N}+1}(\bar{\partial}_{\dot{\mu}}\bar{F}-i\bar{y}_{\dot{\mu}}[\frac{1}{\bar{N}+1}F,\bar{F}]).$ (38)
@@
¯$\ll\bar{\partial}_{\dot{\mu}}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=0,$ (40)
@@
Using this and contracting (32) with $y^{\mu}$ yields
¯$\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\bar{\partial}_{\dot{\mu}}F-i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+2)}F,F],$ (42)
@@
Substituting (37), (38) and (43) into (32) and combining like terms using the gauge algebra Jacobi identity, the final result is
@@
Now we are ready to write down an unfolded system for Yang–Mills theory. Contracting (5) and (44) with the vierbeins yields
@@
$+\frac{3}{2}ey\bar{y}[\frac{1}{(N+1)(N+2)}[\frac{1}{N}F,\bar{F}],F]+ey\bar{y}[\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F],\frac{1}{N}F]$ (55)
@@
A full spectrum of unfolded fields consists of a 1-form $\Omega$ describing Minkowski background, a 1-form of the gauge potential $A$ and the 0-forms of master-fields $F(Y|x)$ and $\bar{F}(Y|x)$ encoding the Yang–Mills tensor together with an infinite tower of its covariant derivatives. Corresponding unfolded equations are (46), (54) and (55) plus a conjugate for $\bar{F}$. The formulation is manifestly diffeomorphism-invariant. 1-forms $\Omega$ and $A$ give rise to two manifest symmetries in accordance with (3): the global (after fixing $\Omega$) Poincaré one and the local Yang–Mills one. The Yang–Mills symmetry is realized as
@@
By construction, the unfolded system is consistent, provided unfolded 0-form fields obey the helicity constraint (15). At the same time, a direct check of (2) seems hardly executable due to the complexity of the equations. One of the possible forms of solution to the system is (11) with primary fields subjected to (4). Let us quickly derive this directly from (55).
@@
Then (11) explicitly realizes the first arrow in (59). The field $\mathcal{F}(Y)$ carries precisely the same information as on-shell $F_{\alpha\alpha}$ does. In a sense, spinors $Y=\{y^{\alpha};\bar{y}^{\dot{\alpha}}\}$ effectively replace space-time coordinates $x^{\alpha\dot{\alpha}}$ for on-shell configurations, hence being conjugate to spinor-helicity variables that resolve light-like momenta $p^{\alpha\dot{\alpha}}=\pi^{\alpha}\bar{\pi}^{\dot{\alpha}}$. This is the way the unfolded system imposes e.o.m. on primary fields: via (59), it maps $4d$ space-time fields onto an effectively $3d$ hypersurface (in the sense that $y^{\alpha}\bar{y}^{\dot{\alpha}}$ is a light-like vector). Note, however, that the role of auxiliary spinors is much more important and sophisticated. As follows directly from (11), they in fact equalize and mix translational and spin degrees of freedom. So they should not be thought of simply as coordinates on some null hypersurface in $4d$ Minkowski space.
@@
In fact, one can start directly from (61) and then make use of (55) in order to fully recover $x$-dependence. It means that although we have derived the unfolded equation (55) starting from postulating (11), for (55) the expression (11) per se is nothing more than just one of many possible forms of solution. In particular, instead of (59) one can think of (55) as defining a $Y$-to-$x$ map (a similar interpretation has been proposed in [25])
@@
[…]
@@
It generates a solution to (55), and hence implicitly to Maxwell equations, for arbitrary $\mathcal{F}_{\alpha\alpha}(y\bar{y})$. This signals that relativistic dynamics in fact can be realized in terms of $Y$ without any reference to a space-time. Corresponding action principle for an arbitrary-mass integer-spin field was constructed in [26] by means of supplementing the set of $Y$ with a Lorentz-invariant proper-time coordinate serving as an evolution parameter.
The unfolding map (62) hints that the unfolded dynamics approach may provide a new way to study the problem of integrability. In particular, the question is whether it is possible to extend (64) to the non-abelian case. If so, an appropriate unfolding map will generate solutions to Yang–Mills equations in Minkowski space.

### v3

@@ Abstract
In this note, we present a novel formulation of $4d$ pure Yang-Mills theory within the unfolded framework of Vasiliev higher-spin gravity. This formulation is first-order and exhibits manifest diffeomorphism and gauge invariance. Our approach builds upon a recently proposed unfolding method, previously applied to scalar electrodynamics. Additionally, we discuss the features of various unfolding maps defined by the unfolded equations.
@@ Introduction
Symmetry considerations serve as one of the most fundamental guiding principles in constructing theories of fundamental interactions. Higher-spin gravity theories provide prime examples, describing systems of interacting massless fields of all spins. To a large extent, these theories are determined by an infinite-dimensional higher-spin gauge symmetry [1]. For a partial overview of the relevant literature, see [2].
Maintaining control over higher-spin symmetry is of fundamental importance. The first example of higher-spin gravity was formulated through the Vasiliev equations [3, 4], a set of classical equations of motion written in the so-called unfolded form. Unfolded equations are first-order differential equations for unfolded fields – which are exterior forms – that possess manifest gauge symmetry. Thus, the unfolded dynamics approach [5, 3, 4, 6, 7, 8] provides a classical first-order formalism that ensures both manifest diffeomorphism and gauge invariance. It also enables effective control over the theory’s degrees of freedom: unfolded fields (which form an infinite spectrum in dynamical field theories) parameterize all d.o.f. of the system. In particular, these properties make the unfolded dynamics approach a promising tool for investigating $AdS/CFT$ correspondence and other dualities [9, 10, 11]. A quantization scheme for classical unfolded theories was put forward in [12] (see also [13, 14, 15, 16] for discussions of quantization of non-Lagrangian field theories).
Beyond higher-spin gravity models [3, 4, 17, 18, 19, 20], relatively few examples of unfolded formulations exist for nonlinear theories [12, 21] (see also [22] for a discussion of nonlinear unfolding in conformal geometry), thereby constraining comprehensive exploration of the approach’s full potential. The primary obstacle was the lack of a practical unfolding algorithm – especially critical for nonlinear theories. Therefore, linear models were mainly studied [23, 24, 25, 26, 27]. In [21], a novel unfolding method was put forward, that enabled the construction of an unfolded formulation of $4d$ scalar electrodynamics. In essence, it involves two steps: first postulating the form of an unfolded master-field, then deriving the corresponding unfolded equations as identities for this master-field.
In this note, we adapt and refine the method of [21] to construct an unfolded formulation of $4d$ on-shell pure Yang–Mills theory.
The note is organized as follows. First, we review basic concepts of the unfolded dynamics approach. Next, we construct the unfolded Yang–Mills equations by fixing the form of the unfolded master-fields and using certain operator relations that we derive. We then discuss the properties of the resulting unfolded system and comment on the unfolding maps it defines. In conclusion, we outline promising directions for future research.
@@
Within the unfolded dynamics approach [5, 3, 4, 6, 7, 8], field theories are formulated via first-order differential equations
@@
on unfolded fields $W^{A}(x)$, which are exterior forms on a space-time manifold $M^{d}$. Here, $A$ collectively denotes all indices of an unfolded field, $\mathrm{d}$ is the exterior derivative on $M^{d}$, and $G^{A}(W)$ is constructed from exterior products of unfolded fields (the wedge symbol is omitted throughout the paper). For each $W^{A}$ there is one and only one unfolded equation (1).
The system (1) must obey the consistency condition
@@
which follows from $\mathrm{d}^{2}\equiv 0$ and constrains possible forms of $G^{A}$. Considering this, the unfolded equations (1) possess manifest gauge symmetry [^1] Rigorously, this holds provided the consistency is dimension-independent – a property satisfied by all known examples [8].
@@
with each $(n>0)$-form field $W^{A}$ giving rise to a gauge symmetry with (n1)(n-1)-form parameter $\varepsilon^{A}(x)$, while 0-form fields transform only under gauge symmetries of 1-form fields through the second term in (3).
In dynamical field theories, the spectrum of unfolded fields is infinite, as they parameterize all physical d.o.f. Typically, the space of unfolded fields admits a grading bounded from below, so that equations (1) express (perhaps, in a very complicated nonlinear way) higher-grade fields through derivatives of lower-grade ones. In addition, (1) may implicitly impose dynamical constraints on the lowest-grade fields, rendering the system on-shell. Equations (1) define what is known as a free differential algebra [28] (see also [29, 16] for discussions of the relation between the unfolded framework and various mathematical structures).
Ultimately, an unfolded system (1) describes a theory of lowest-grade fields (also referred to as primaries), possibly subject to some differential constraints (e.o.m. or whatever), while higher-grade fields constitute (infinite) towers of their covariant differential descendants. Crucially, the unfolded formulation maintains manifest diffeomorphism and gauge invariance and provides complete control over d.o.f. of the theory. These features establish unfolded dynamics approach as a powerful framework for investigating higher-spin gravity. However, these advantages extend beyond higher spins, offering potential applications to conventional field theories. In this note, our goal is to reformulate $4d$ pure Yang–Mills theory in the unfolded form (1), applying the unfolding method introduced in [21].
@@
In $sl(2,\mathbb{C})$-spinor notation, the $4d$ Yang–Mills equations together with Bianchi identities are
@@
where the (anti-)self-dual components of the field strength tensor are
$F_{\alpha\alpha}:=\frac{\partial}{\partial x^{\alpha\dot{\beta}}}A_{\alpha}{}^{\dot{\beta}}-i[A_{\alpha\dot{\beta}},A_{\alpha}{}^{\dot{\beta}}],\hskip 10.00002pt\bar{F}_{\dot{\alpha}\dot{\alpha}}:=\frac{\partial}{\partial x^{\beta\dot{\alpha}}}A^{\beta}{}_{\dot{\alpha}}-i[A_{\beta\dot{\alpha}},A^{\beta}{}_{\dot{\alpha}}$], (5)
@@
satisfying
$\mathrm{D}_{\alpha\dot{\alpha}},\mathrm{D}_{\beta\dot{\beta}}]=-i\epsilon_{\alpha\beta}\bar{F}_{\dot{\alpha}\dot{\beta}}-i\epsilon_{\dot{\alpha}\dot{\beta}}F_{\alpha\beta}.$ (7)
The antisymmetric spinor metric
@@
raises/lowers spinor indices via
@@
For multispinors, indices denoted with the same letter are either contracted or symmetrized, depending on their relative positions, (similarly for dotted indices)
@@
In order to unfold Yang–Mills theory, one has to introduce, on top of primaries $F_{\alpha\alpha}$ and $\bar{F}_{\dot{\alpha}\dot{\alpha}}$, the infinite towers of all their differential on-shell descendants. This can be conveniently performed by means of auxiliary commuting spinors [^2] In $4d$ Vasiliev equations [3, 4], these spinors become generators of an associative higher-spin gauge algebra through a specific non-commutative star product defined on them. $Y=(y^{\alpha},\bar{y}^{\dot{\alpha}})$, then the whole towers get packed into unfolded Yang–Mills master-fields, which we postulate to be of the form
@@
together with an infinite sequence of their fully symmetrized traceless covariant derivatives of all orders. This constitutes the set of all independent covariant descendants of the primary Yang–Mills tensor, since antisymmetrizations and contractions are determined by (4) and (7). By construction, unfolded master-fields (11) inherit the adjoint representation of the gauge algebra of primary tensors.
@@
where $\partial_{\alpha}$ and $\bar{\partial}_{\dot{\alpha}}$ are $y^{\alpha}$- and $\bar{y}^{\dot{\alpha}}$-derivatives. From their definitions (11), master-fields obey
@@
which corresponds to helicities $\pm 1$ (in higher-spin gravity, strength tensors with $|N-\bar{N}|F=2sF$ describe spin-$s$ massless fields) and, further,
@@
so that the primary has 0-grade, while $n$-grade unfolded fields represent its covariant derivatives of $n$-th order, as follows from (11).
To formulate the unfolded equations (1), one needs to express the derivatives of the unfolded fields in algebraic terms. In our particular case, the task is to express $\mathrm{D}_{\alpha\dot{\beta}}F$ for $F$ (11) in terms of $Y$ and $\partial/\partial Y$ acting on $F$ and $\bar{F}$ (and the same for $\mathrm{D}_{\alpha\dot{\beta}}\bar{F}$). Before we start processing (11), let us work out some general operator relations that simplify the analysis implied by the method of [21]. Combining a commutator formula
$\hat{\mathrm{A}},e^{\hat{\mathrm{D}}}]=\int_{0}^{1}dte^{t\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-t\hat{\mathrm{D}}}e^{\hat{\mathrm{D}}}$ (19)
with an Euler-operator representation of a homotopy integral
$\int_{0}^{1}dtt^{k}F(tz)=\frac{1}{z\frac{\partial}{\partial z}+1+k}F(z),$ (20)
one obtains
@@
where the inverse Euler operator is understood (when $F(0)=0$) as
@@
Applying this formula to an operator of the form
@@
$\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}}e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}$. (24)
@@
[…]
@@
$\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}(N_{\hat{\mathrm{D}}}-1)}e^{\hat{\mathrm{D}}}[\hat{\mathrm{D}},[\hat{\mathrm{A}},\hat{\mathrm{D}}]]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+[[\hat{\mathrm{A}},\hat{\mathrm{D}}],\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}$. (26)
For convenience, we denote an "unfolding map" for an arbitrary field $\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)$ taking values in the adjoint representation as
@@
Then from (26) one has (here $C\equiv\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)$ )
$\ll\partial_{\mu}\mathrm{C}\gg=(\partial_{\mu}-\mathrm{D}_{\mu\dot{\mu}}\bar{y}^{\dot{\mu}})\ll\mathrm{C}\gg+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\ll\mathrm{C}\gg],$ (29)
$\ll\bar{\partial}_{\dot{\mu}}\mathrm{C}\gg=(\bar{\partial}_{\dot{\mu}}-\mathrm{D}_{\mu\dot{\mu}}y^{\mu})\ll\mathrm{C}\gg+i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+1)(\bar{N}+2)}F,\ll\mathrm{C}\gg].$ (30)
From here on, all derivatives inside the angle brackets always act only on an expression within brackets and never differentiate unfolding exponents of (27). The square brackets stand for the commutator in the gauge Lie algebra and the Euler operators of (26) are expressed in terms of the spinor Euler operators (14). Master-fields $F$ and $\bar{F}$ arise in (29)-(30) through (7) without assuming Yang–Mills equations (4).
Now, let us get down to solving the problem. Direct application of (24) to $\mathrm{D}_{\mu\dot{\mu}}F(Y|x)$ gives
@@
The task is to eliminate all covariant derivatives and angle brackets on the r.h.s. by re-expressing them in terms of unfolded master-fields $F$ and $\bar{F}$ acted on by $Y$’s and $\frac{\partial}{\partial Y}$’s. To accomplish this, we have the following tools at our disposal: Yang–Mills equations (4) together with (7), the relations (29) and (30), the Schouten identities for spinors and the Jacobi identity of the gauge Lie algebra.
@@
Therefore, one needs to process $\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F$. Contracting (32) with $\bar{y}^{\dot{\mu}}$ gives
@@
$\ll\bar{\partial}_{\dot{\mu}}\bar{F}_{\dot{\alpha}\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg=\frac{2}{\bar{N}+1}(\bar{\partial}_{\dot{\mu}}\bar{F}-i\bar{y}_{\dot{\mu}}[\frac{1}{\bar{N}+1}F,\bar{F}]).$ (38)
@@
$\ll\bar{\partial}_{\dot{\mu}}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=0,$ (40)
@@
Using this result and contracting (32) with $y^{\mu}$ yields
$\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\bar{\partial}_{\dot{\mu}}F-i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+2)}F,F],$ (42)
@@
Substituting (37), (38) and (43) into (32) and combining like terms using the gauge algebra Jacobi identity, one finally expresses the covariant derivative of $F$ in $Y$-terms
@@
Now we are ready to write down an unfolded system for Yang–Mills theory, which is the main result of the note. Contracting (5) and (44) with the vierbeins yields
@@
$+\frac{3}{2}ey\bar{y}[\frac{1}{(N+1)(N+2)}[\frac{1}{N}F,\bar{F}],F]+ey\bar{y}[\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F],\frac{1}{N}F],$ (55)
@@
A full spectrum of unfolded fields consists of a 1-form $\Omega$ describing Minkowski background, a 1-form of the gauge potential $A$ and the 0-forms of master-fields $F(Y|x)$ and $\bar{F}(Y|x)$ encoding the Yang–Mills tensor together with an infinite tower of its covariant derivatives. The corresponding unfolded equations are (46), (54) and (55) plus a conjugate for $\bar{F}$. The formulation is manifestly diffeomorphism-invariant. 1-forms $\Omega$ and $A$ give rise to two manifest symmetries in accordance with (3): the global (after fixing $\Omega$) Poincaré one and the local Yang–Mills one. The Yang–Mills symmetry is realized as
@@
By construction, the unfolded system is consistent, provided unfolded 0-form fields obey the helicity constraint (15). At the same time, a direct check of (2) seems hardly executable due to the complexity of the equations. Note that the system is at most cubic in master-field 0-forms, which is surprising on its own. One of the possible forms of solution to the system is (11) with primary fields subjected to (4). Let us quickly derive this directly from (55).
@@
Then (11) explicitly realizes the first arrow in (59). The field $\mathcal{F}(Y)$ carries precisely the same information as on-shell $F_{\alpha\alpha}$ does. In a sense, spinors $Y=\{y^{\alpha};\bar{y}^{\dot{\alpha}}\}$ effectively replace space-time coordinates $x^{\alpha\dot{\alpha}}$ for on-shell configurations, hence being conjugate to spinor-helicity variables that resolve light-like momenta $p^{\alpha\dot{\alpha}}=\pi^{\alpha}\bar{\pi}^{\dot{\alpha}}$. This is the way the unfolded system imposes e.o.m. on primary fields: via (59), it maps $4d$ space-time fields onto an effectively $3d$ hypersurface (in the sense that $y^{\alpha}\bar{y}^{\dot{\alpha}}$ is a light-like vector). Note however, that the role of auxiliary spinors is much more important and sophisticated. As follows directly from (11), they in fact equalize and mix translational and spin degrees of freedom. So they should not be thought of simply as coordinates on some null hypersurface in $4d$ Minkowski space.
@@
In fact, one can start directly from (61) and then make use of (55) in order to fully recover $x$-dependence. This implies that although we have derived the unfolded equation (55) starting from postulating (11), for (55) the expression (11) per se is nothing more than just one of many possible forms of solution. In particular, instead of (59) one can think of (55) as defining a $Y$-to-$x$ map (a similar interpretation has been proposed in [31])
@@
$(N-\bar{N})\mathcal{F}(Y)=2\mathcal{F}(Y)\hskip 10.00002pt\Rightarrow\hskip 10.00002pt\mathcal{F}(Y)=\mathcal{F}_{\alpha\alpha}(y\bar{y})y^{\alpha}y^{\alpha}.$ (63)
@@
It generates a solution to (55), and hence implicitly to Maxwell equations, for arbitrary $\mathcal{F}_{\alpha\alpha}(y\bar{y})$. This indicates that relativistic dynamics in fact can be realized in terms of $Y$ without any reference to a space-time. The corresponding action principle for an arbitrary-mass integer-spin field was constructed in [32] by means of supplementing the set of $Y$ with a Lorentz-invariant proper-time coordinate serving as an evolution parameter.
From this perspective, Yang–Mills theory is characterized by the unconstrained field $\mathcal{F}(Y)=\mathcal{F}_{\alpha\alpha}(y\bar{y})y^{\alpha}y^{\alpha}$ and its conjugate $\bar{\mathcal{F}}$. These field form representations of both the gauge and Poincaré algebras, as defined by the unfolded system via the general formula (3). At first glance, this might suggest the theory is already solved, since no equations constrain $\mathcal{F}$ and $\bar{\mathcal{F}}$. However, the key complication preserving the theory’s non-triviality lies in the nonlinear realization of Poincaré translations in $\mathcal{F}$ and $\bar{\mathcal{F}}$, as evident from (55). Considering this, the unfolding map (62) suggests that the unfolded dynamics approach may provide a new way to study the problem of integrability of classical e.o.m. and constructing exact solutions. In particular, the question is whether it is possible to extend (64) to the non-abelian case. If so, an appropriate unfolding map will generate solutions to Yang–Mills equations in Minkowski space. Additionally, the unfolded framework provides a systematic method for identifying theory invariants: all gauge-invariant conserved charges are classified through the cohomology of a certain operator determined by the unfolded equations [8].